If the nuclear radius of Al7 is 3.6 fm, the approximate nuclear radius of Cu64
in fermi is
1
Expert's answer
2021-10-04T17:36:49-0400

"r=r_0A^{\frac 13},"

"\frac{r_{Cu}}{r_{Al}}=(\frac{{A_{Cu}}}{A_{Al}} )^{\frac 13}=(\frac{64}{27})^{\frac 13}=\frac 43,"

"r_{Cu}=\frac 43\cdot 3.6=4.8~\text{fm}."